# Requirements Coverage-Guided Minimization for Natural Language Test Cases

### v1

@@ 6. Threats to Validity
External Validity. The generalizability of our findings may be limited due to the specific domain and nature of the test suites provided by our industry partner. The performance of our approach might differ when applied to other domains or software systems with different requirement structures, test suite characteristics, or fault profiles. To mitigate this threat, we selected test suites that vary in size and redundancy levels to evaluate the robustness of our approach under diverse configurations. Furthermore, our approach is designed to be domain-agnostic, relying solely on textual information from requirements and test cases, which facilitates potential adaptation to other domains. Last, our industry partner is following standard development practices in the automotive domain, imposed by standards, that are similar to other safety-critical domains where functional safety is important. Future work includes evaluating our approach on publicly available datasets from different domains to further assess the generalizability.
@@
Another threat to external validity arises from the choice of embedding techniques that is limited by the data privacy constraints of our industry partner. This limitation may have affected the overall performance of our approach. Nevertheless, to mitigate this threat, we evaluated seven different embedding methods within the bounds of our available computational resources. This broad comparison helps reduce concerns regarding internal validity, despite the constraints on embedding selection.
Conclusion Threats. To reduce the impact of randomness and increase the reliability of our results, we run each experiment ten times with different random seeds and report the average performance. This mitigates the risk of drawing conclusions based on outlier results or fluctuations due to stochastic components in the minimization process.
@@ 7. Related Work
A number of approaches have been proposed to support TSM, which can be classified in to three categories (Khan et al., 2018; Cruciani et al., 2019; Pan et al., 2023): greedy (Noemmer and Haas, 2019; Miranda and Bertolino, 2017), clustering (Coviello et al., 2018; Liu et al., 2011; Viggiato et al., 2022; Cruciani et al., 2019), and search-based (Zhang et al., 2019; Hemmati et al., 2013; Pan et al., 2023, 2024; Marchetto et al., 2017), plus hybrid combinations (Xia et al., 2021; Anwar et al., 2019; Yoo and Harman, 2010) thereof.
@@
Greedy-based approaches follow a heuristic, step-by-step selection strategy, where the test case that provides the most benefit toward the objective is chosen at each iteration, until the given constraints (e.g., full requirement coverage or code path coverage) are satisfied. Miranda et al. (Miranda and Bertolino, 2017) and Noemmer et al. (Noemmer and Haas, 2019) both employed greedy heuristics for TSM based on statement coverage. Their approaches iteratively selected test cases to maximize code coverage, achieving reasonable FDRs while significantly reducing the number of executed test cases. Greedy-based approaches are computationally efficient and easy to implement but may suffer from suboptimal global performance due to their locally optimal decisions
@@
Clustering-based approaches group similar test cases based on syntactic or semantic similarities, such as those derived from natural language descriptions or execution behavior. The goal is to reduce redundancy by selecting representative test cases from each cluster. These approaches are particularly useful in settings where test cases are written in natural language and code coverage information is unavailable. Viggiato et al. (Viggiato et al., 2022) proposed a comprehensive approach that combines text embedding, similarity measurement, and clustering to identify and remove similar test cases written in natural language, where they split requirement test cases into test steps and then calculate the similarity between test steps. Cruciani et al. (Cruciani et al., 2019) proposed FAST-R, a black-box TSM approach that leverages the source code of test cases without requiring execution or coverage information. FAST-R transforms test code into vector representations using a term frequency model, followed by dimensionality reduction via random projection. Similarly, Coviello et al. (Coviello et al., 2018) proposed a clustering-based approach using Hierarchical Agglomerative Clustering to perform TSM.
@@
Search-based approaches formulate the minimization problem as an optimization task and utilize metaheuristic algorithms (e.g., genetic algorithms, simulated annealing) to explore the space of possible test subsets. These methods are capable of balancing multiple conflicting objectives, such as minimizing the number of test cases while maximizing fault detection capability and maintaining requirement coverage. Hemmati et al. (Hemmati et al., 2013) applied a search algorithm that minimizes test suites based on model-derived test case similarity. Similarly, Zhang et al.(Zhang et al., 2019) proposed UncerTest, a model-based test case minimization framework that leverages multi-objective search algorithms, including NSGA-II, MOCell(Nebro et al., 2009), and SPEA2 (Zitzler et al., 2001). ATM (Pan et al., 2023) is a black-box TSM approach that combines test code similarity analysis with evolutionary search. It represents test code using Abstract Syntax Trees and employs Genetic Algorithm and Non-Dominated Sorting Genetic Algorithm II (Meyarivan et al., 2002) as the underlying evolutionary algorithms. Furthermore, LTM (Pan et al., 2024) leverages LLMs (CodeBERT, UniXcoder, and CodeLlama) to compute similarity between test cases and employs a Genetic Algorithm to search for an optimal subset of test cases under a fixed budget. MORE+ (Marchetto et al., 2017) is a multi-objective TSM approach that uses NSGA-II to optimize structural, functional, and cost-related objectives, aiming to enhance fault detection while reducing redundancy and execution time. Although typically more computationally intensive than greedy and clustering approaches, search-based techniques support multiple minimization objectives and often achieve better performance for TSM.
@@
Hybrid approaches integrate complementary methods into a unified framework to balance multiple objectives, including reducing test suite size, preserving fault detection capability, and satisfying requirement or coverage constraints. Xia et al. (Xia et al., 2021) applies clustering-based test suite reduction combined with evolutionary multi-objective optimization. Anwar et al. (Anwar et al., 2019) combined genetic algorithm and particle swarm optimization algorithm to optimize regression test suites. Yoo et al. (Yoo and Harman, 2010) employs a hybrid multi-objective genetic algorithm that integrates the efficient approximation of the greedy approach with the capability of population-based genetic algorithms to generate higher-quality Pareto fronts.
@@
While promising, existing techniques primarily focus on code-based solutions for minimization. In contrast, our work addresses minimization in the context of requirement-driven testing where test cases are specified in natural language and are thus introducing distinct challenges. Moreover, current approaches do not explicitly support TSM under a fixed budget while simultaneously enforcing full requirement coverage as a hard constraint. To the best of our knowledge, this is also the first work to investigate the impact of redundancy levels on fault detection capability under such constraints, providing valuable insights for future research and practical applications. Note that we retained FAST-R as one of the baselines though it cannot simultaneously satisfy both the fixed-size minimization budget and the full coverage adequacy constraint. This is because FAST-R can be applied in two scenarios (i.e., fixed-size budget and adequate), which enables experimental comparison with our approach.
@@ 8. Conclusion
In this paper, we presented RTM, a novel and practical approach for test suite minimization guided by requirements coverage. As it is often the case in domains where functional safety is crucial and must be demonstrated through adherence to standards, test cases are specified in natural language (test steps), are traced to requirements and are implemented manually.
By combining advanced text embedding techniques, multiple similarity metrics, and a Genetic Algorithm (GA) initialized using multiple strategies, RTM effectively reduces test suite size while strictly preserving full requirement coverage under a fixed budget constraint. Its evaluation on an industrial automotive dataset demonstrates its superior performance in Fault Detection Rate (FDR) across various minimization budgets. In comparison with multiple baselines, RTM consistently achieved significantly higher FDR, confirming its effectiveness in an industrial setting. Furthermore, our exploration of test suite redundancy levels revealed that higher redundancy significantly improves FDR outcomes, while higher redundancy combined with low minimization budgets tends to significantly increase the risk of undetected faults, thus highlighting the importance of considering redundancy characteristics in test suite minimization.
In conclusion, our results establish RTM as a scalable and reliable minimization solution for requirement-based testing (validation) in safety-critical systems. In the future, we plan to apply RTM to a broader range of application scenarios to further evaluate and enhance its generalizability.
@@ Acknowledgements.
This work was supported by a research grant from Aptiv, as well as by the Discovery Grant and Canada Research Chair programs of the Natural Sciences and Engineering Research Council of Canada (NSERC) and Research Ireland grant 13/RC/2094-2.

### v2

@@ 6. Threats to Validity
External Validity. The generalizability of our findings may be limited due to the specific domain and nature of the test suites provided by our industry partner. The performance of our approach might differ when applied to other domains or software systems with different requirement structures, test suite characteristics, or fault profiles. To mitigate this threat, we selected test suites of varying sizes and redundancy levels to evaluate the robustness of our approach across diverse configurations. Furthermore, our approach is domain-agnostic, relying solely on textual information from requirements and test cases, thereby facilitating adaptation to other domains. Lastly, our industry partner follows standard development practices in the automotive domain, as defined by standards similar to those in other safety-critical domains where functional safety is important. Future work includes evaluating our approach on publicly available datasets from different domains to further assess the generalizability.
@@
Another threat to external validity arises from the limited choice of embedding techniques, constrained by the data privacy requirements of our industry partner. This limitation may have affected the overall performance of our approach. Nevertheless, to mitigate this threat, we evaluated seven different embedding methods within the bounds of our available computational resources. This broad comparison helps reduce concerns regarding internal validity, despite the constraints on embedding selection.
Construct Threats. A potential construct threat concerns the maintenance cost of TF-IDF representations when new test cases are added. In our context, this risk is limited because the dataset originates from automotive system requirement-based test cases, which follow well-defined structural templates and rely on a stable, domain-specific vocabulary. As a result, the introduction of new terms is infrequent and typically occurs only during major version updates, making the need for corpus-wide re-computation of TF-IDF vectors relatively unlikely. Moreover, as discussed in Section 5.4, the preparation time of TF-IDF is scalable for large test suites, and the maintenance cost is therefore acceptable. However, in domains where the vocabulary evolves rapidly or new linguistic expressions appear frequently (e.g., user-generated content or dynamic project documentation), we recommend using domain-specific or LLM-based embeddings that can better adapt to incremental updates without requiring full re-vectorization.
Conclusion Threats. To reduce the impact of randomness and increase the reliability of our results, we run each experiment ten times with different random seeds and report the average performance. This mitigates the risk of drawing conclusions based on outlier results or fluctuations arising from stochastic components in the minimization process.
@@ 7. Related Work
Several approaches have been proposed to support TSM, which can be classified in to three categories (khan2018systematic; cruciani2019scalable; pan2023atm): greedy (noemmer2019evaluation; miranda2017scope), clustering (coviello2018clustering; liu2011user; viggiato2022identifying; cruciani2019scalable), and search-based (zhang2019uncertainty; hemmati2013achieving; pan2023atm; DBLP:journals/tse/PanGB24; marchetto2017combining), plus hybrid combinations (xia2021test; anwar2019hybrid; yoo2010using) thereof.
@@
Greedy-based approaches follow a heuristic, step-by-step selection strategy, where the test case that provides the most benefit toward the objective is chosen at each iteration, until the given constraints (e.g., full requirement coverage or code path coverage) are satisfied. Miranda et al. (miranda2017scope) and Noemmer et al. (noemmer2019evaluation) both employed greedy heuristics for TSM based on statement coverage. Their approaches iteratively selected test cases to maximize code coverage, achieving reasonable FDRs while significantly reducing the number of executed test cases. Greedy-based approaches are computationally efficient and easy to implement, but may suffer from suboptimal global performance due to their locally optimal decisions. Moreover, given that, in our dataset, each test case covers only one requirement, the coverage-based greedy strategy–iteratively selecting the test cases that maximize the coverage until the full coverage constraint is satisfied–is equivalent to Random Minimization with the requirement constraint (RM-Req), as described in Section 4.4
@@
Clustering-based approaches group similar test cases based on syntactic or semantic similarities, such as those derived from natural language descriptions or execution behavior. The goal is to reduce redundancy by selecting representative test cases from each cluster. These approaches are particularly useful in settings where test cases are written in natural language and code coverage information is unavailable. Viggiato et al. (viggiato2022identifying) proposed a comprehensive approach that combines text embedding, similarity measurement, and clustering to identify and remove similar test cases written in natural language. They split the requirement test cases into test steps and then calculate the similarity between test steps. Cruciani et al. (cruciani2019scalable) proposed FAST-R, a black-box TSM approach that leverages the source code of test cases without requiring execution or coverage information. FAST-R transforms test code into vector representations using a term frequency model, followed by dimensionality reduction via random projection. Similarly, Coviello et al. (coviello2018clustering) proposed a clustering-based approach using Hierarchical Agglomerative Clustering to perform TSM.
@@
Search-based approaches formulate the minimization problem as an optimization task and utilize metaheuristic algorithms (e.g., GA, simulated annealing) to explore the space of possible test subsets. These methods can balance multiple conflicting objectives, such as minimizing the number of test cases while maximizing fault detection capability and maintaining requirement coverage. Hemmati et al. (hemmati2013achieving) applied a search algorithm that minimizes test suites based on model-derived test case similarity. Similarly, Zhang et al.(zhang2019uncertainty) proposed UncerTest, a model-based test case minimization framework that leverages multi-objective search algorithms, including NSGA-II, MOCell(nebro2009mocell), and SPEA2 (zitzler2001spea2). ATM (pan2023atm) is a black-box TSM approach that combines test code similarity analysis with evolutionary search. It represents test code using Abstract Syntax Trees and employs GA and Non-Dominated Sorting Genetic Algorithm II (meyarivan2002fast) as the underlying evolutionary algorithms. Furthermore, LTM (DBLP:journals/tse/PanGB24) leverages LLMs (CodeBERT, UniXcoder, and CodeLlama) to compute similarity between test cases and employs a GA to search for an optimal subset of test cases under a fixed budget. MORE+ (marchetto2017combining) is a multi-objective TSM approach that uses NSGA-II to optimize structural, functional, and cost-related objectives, aiming to enhance fault detection while reducing redundancy and execution time. Although typically more computationally intensive than greedy and clustering approaches, search-based techniques support multiple minimization objectives and often achieve better performance for TSM.
@@
Hybrid approaches integrate complementary methods into a unified framework to balance multiple objectives, including reducing test suite size, preserving fault detection capability, and satisfying requirement or coverage constraints. Xia et al. (xia2021test) apply clustering-based test suite reduction combined with evolutionary multi-objective optimization. Anwar et al. (anwar2019hybrid) combined a GA and particle swarm optimization to optimize regression test suites. Yoo et al. (yoo2010using) employ a hybrid multi-objective GA that integrates the efficient approximation of the greedy approach with the population-based GA’s ability to generate higher-quality Pareto fronts.
@@
While promising, existing techniques primarily focus on code-based solutions for minimization. In contrast, our work addresses minimization in the context of requirement-driven testing, where test cases are specified in natural language, which introduces distinct challenges. Moreover, current approaches do not explicitly support TSM under a fixed budget while simultaneously enforcing full requirement coverage as a hard constraint. To this end, we adopt a GA, as the SOTA TSM work (i.e., ATM and LTM), and augment it with an explicit requirement coverage constraint that guarantees $100\%$ coverage under a fixed minimization budget. Note that we did not use NSGA-II, as reported in the ATM work, since it achieves comparable performance while being substantially more time-consuming than the GA. To the best of our knowledge, this is also the first work to investigate the impact of redundancy levels on fault detection capability under such constraints, providing valuable insights for future research and practical applications. Note that we retained FAST-R as one of the baselines, though it cannot simultaneously satisfy both the fixed-size minimization budget and the full coverage adequacy constraint. This is because FAST-R can be applied in two scenarios (i.e., fixed-size budget and adequate), which enables experimental comparison with our approach.
@@ 8. Conclusion
In this paper, we presented RTM, a novel and practical approach for test suite minimization guided by requirements coverage. As is often the case in domains where functional safety is crucial and must be demonstrated through adherence to standards, test cases are specified in natural language (test steps), traced to requirements, and implemented manually.
RTM combines the text embedding technique, distance function, and a GA to effectively reduce test suite size while strictly preserving full requirement coverage under a fixed budget constraint. We investigate three preprocessing methods for test cases, seven text embedding techniques, three distance functions, and three initialization strategies for the GA. Its evaluation on an industrial automotive dataset demonstrates its superior performance in Fault Detection Rate (FDR) across various minimization budgets. Compared with multiple baselines, while being efficient in runtime, RTM achieves higher FDR across most minimization budgets, making it a better choice in many practical contexts. Furthermore, our exploration of test suite redundancy levels revealed that higher redundancy significantly improves FDR, while higher redundancy combined with low minimization budgets tends to significantly increase the risk of undetected faults, highlighting the importance of considering redundancy characteristics in test suite minimization.
In conclusion, our results establish RTM as a scalable and reliable minimization solution for requirement-based testing (validation) in safety-critical systems. In the future, we plan to apply RTM to a broader range of application scenarios to evaluate further and enhance its generalizability.
@@ Acknowledgements.
This work was supported by a research grant from Aptiv, the Discovery Grant and Canada Research Chair programs of the Natural Sciences and Engineering Research Council of Canada (NSERC), and a Research Ireland grant 13/RC/2094-2.